### v2

# Implicit Personalization in Language Models: A Systematic Study

Abstract
Implicit Personalization (IP) is a phenomenon of language models inferring a user’s background from the implicit cues in the input prompts and tailoring the response based on this inference. While previous work has touched upon various instances of this problem, there lacks a unified framework to study this behavior. This work systematically studies IP through a rigorous mathematical formulation, a multi-perspective moral reasoning framework, and a set of case studies. Our theoretical foundation for IP relies on a structural causal model and introduces a novel method, indirect intervention, to estimate the causal effect of a mediator variable that cannot be directly intervened upon. Beyond the technical approach, we also introduce a set of moral reasoning principles based on three schools of moral philosophy to study when IP may or may not be ethically appropriate. Equipped with both mathematical and ethical insights, we present three diverse case studies illustrating the varied nature of the IP problem and offer recommendations for future research. [^1] Our code is at https://github.com/jiarui-liu/IP, and our data is at https://huggingface.co/datasets/Jerry999/ImplicitPersonalizationData.
1 Introduction
Let’s begin with a brain teaser: What color is a football? As illustrated in Figure 1, we first infer from the spelling “color” – as opposed to “colour” – that the user speaks American English. Therefore, we answer “Brown,” in contrast to the black and white pattern typically for a football in British English.
Figure 1: Overview of the general formulation of IP, where the model infers the user background from the text input, and then customizes the response.
Inspired by this example, we propose the concept of Implicit Personalization (IP). Grounded in a structural causal model (SCM; Peters et al., 2017; Pearl, 2009), we define IP as a process that first infers a user’s background from the way a question is posed, and then tailors the response to fit this background, as in Figure 1. While many studies have separately explored different aspects of this problem (Flek, 2020; Raharjana et al., 2021; Eloundou et al., 2024), we still lack a community-wide standardized framework to study these phenomena. The absence of a common framework leads to divergent perspectives: some studies view it positively, suggesting that incorporating inferred user demographics can enhance NLP performance by personalized responses (Hovy, 2015; Benton et al., 2016; Sasaki et al., 2018; Salemi et al., 2024; Chen et al., 2023), whereas others criticize it negatively for introducing biases in model responses towards underrepresented groups (Bolukbasi et al., 2016; Garg et al., 2018; Arora et al., 2023; Das et al., 2023; He et al., 2024; Kantharuban et al., 2024), or for fostering flattery to satisfy users regardless of the accuracy of the information provided (Sharma et al., 2023a; Wei et al., 2023; Wang et al., 2023).
To this end, we point out that despite the varying terminologies and opinions, all these works fundamentally deal with an instance of IP. Focusing on the essence, our work systematically analyzes IP, by proposing several key research questions and providing answers to them:
RQ1: What is IP? – for which we provide a rigorous mathematical formalization (Section 2). RQ2: How to detect IP in large language models (LLMs)? – for which we provide a rigorous mathematical formulation and a set of case studies (Sections 2, 4, 5, 6 and 7). RQ3: What are the moral implications of IP? – for which we propose a framework for ethical engagement (Section 3). RQ4: How to improve future models to ensure their IP behavior aligns with ethical standards? – for which we provide a list of suggestions for developers and the community (Section 8).
By answering the above questions, our work contributes a “full-stack” systematic study on IP: In the mathematical framework, we ground IP in an SCM Peters et al. (2017); Pearl (2009), and then propose an indirect intervention method to test the causal effect in the LLM-specific, diamond-shaped causal graph in Figure 1 with un-intervenable mediators (see technical deductions in Section 2). After the technical formulation of IP, we provide a moral reasoning framework (Section 3), which connects the ethical considerations of IP to major schools of moral philosophy, including consequentialism Mill (2016); Parfit (1987), deontology Kant and Schneewind (2002); Ross (2002), and contractualism Rawls (2017); Scanlon (2000).
To illustrate the usefulness of our theoretical formulations, we present three diverse case studies that feature different instances of the IP problem (Section 4): (1) cultural adaptation, where IP is a desired model behavior (Section 5), (2) education disparity, where IP is unethical (Section 6), and (3) echo chamber, which has mixed implications for IP (Section 7). Finally, we conclude with recommendations for future research and an outlook for the community (Section 8).
2 A Mathematical Framework for IP
2.1 Causal Graph Formulation
In general, any NLP system has the functional behavior $f:\bm{x}\mapsto\bm{y}$, where the user inputs the text $\bm{x}$, and the model generates a response $\bm{y}$, as in Figure 1. Formally, IP is a sub-process within this functional behavior. In our running example, where $\bm{x}=$“What color is a football?”, the model response $\bm{y}$ should be “Brown” if IP takes place, whereas the general answer would mention both possibilities, e.g., “An American football is Brown, and a soccer ball is usually black and white.”
To evaluate the existence of IP, we model the response generation process of IP with a causal graph $\mathcal{G}$ in Figure 1. In general, a causal graph $\mathcal{G}=(\bm{V},\bm{E})$ is a good mathematical formulation for modeling cause-and-effect relationships among a set of random variables $\bm{V}$, where each edge $e_{ij}\in\bm{E}$ indicates whether the $i$-th random variable $V_{i}$ is a direct cause for the $j$-th random variable $V_{j}$ Pearl (2009); Peters et al. (2017).
In the formulated causal graph of this study in Figure 1, the language model parses $\bm{x}$, extracting its semantics $\bm{s}$—for instance, querying the color of a ball in football—and may infer the user’s background $b$ from a set of categories $\mathcal{B}$ (e.g., American English speaker). The response $\bm{y}$ is generated based on $\bm{s}$, optionally tailored to the user’s background $b$ using IP, or focusing solely on the semantics without customization.
2.2 Problem Statement
Based on our causal graph we introduce a mathematical formulation for implicit personalization. The key research question, namely “Does IP take place in LLMs?”, can be formulated as a question of causal inference, namely “Does user background ${B}$ have a causal effect on the LLM response $\bm{Y}$?”
If the cause $B$ is binary, and the effect variable $Y$ is a real number $Y\in\mathbb{R}$, this question is usually handled by estimating the average treatment effect (ATE) Pearl (1995); Pearl et al. (2000):
$\quad\mathrm{ATE}(B\rightarrow Y)$ (1)
$:=\mathbb{E}[Y|\operatorname*{do}(B=1)]-\mathbb{E}[Y|\operatorname*{do}(B=0)]~{},$ (2)
where $\operatorname*{do}(\cdot)$ denotes an intervention on the variable. The overall estimand calculates the expected change in $Y$ caused by switching $B$ from 0 to 1.
2.2.0.0.1 Quantifying the Interventional Effect: Moving from Numerical Changes to Text Changes
However, the challenge in our study is that the effect variable $\bm{Y}$ is essentially the response of an LLM in natural language, so we can no longer simplify the question to an averaged single value, but need to take into account the entire distribution $P(\bm{Y}|\operatorname*{do}(B=1))$ versus $P(\bm{Y}|\operatorname*{do}(B=0))$ under an intervention on $B$. Using the distributional change to model the causal influence was also discussed in Janzing et al. (2013) in a statistical setting. The general definition of the existence of a causal effect is as follows:
Definition 1
In the given causal graph $\mathcal{G}$, there is a causal effect from $B$ to $\mathbf{Y}$ if there exist $b_{0},b_{1}\in\mathcal{B}$, such that
$P_{\bm{Y}}^{\operatorname*{do}(B=b_{0})}\neq P_{\bm{Y}}^{\operatorname*{do}(B=b_{1})}.$ (3)
The intuition behind this is we first intervene on $B$ by setting it to different values $b_{0}$ and $b_{1}$, and then compare whether the intervened probability distributions of $\bm{Y}$ are identical. If not, then it implies that the LLM performs IP to generate different responses $\bm{Y}$ for different backgrounds $B$. Hence, our research question becomes identifying whether the pair-wise distributional change of $\bm{Y}$ is significant when perturbing $B$. We can perform the following deduction:
IP takes place in the LLM (4)
[…]
 $&\exists b_{0},b_{1}\in\mathcal{B},\text{s.t. }P_{\bm{Y}}^{\operatorname*{do}(B=b_{0})}\neq P_{\bm{Y}}^{\operatorname*{do}(B=b_{1})}.\end{split}$ (6)
The deduction proposed in Eq. 6 can be evaluated using a paired-samples test while controlling the semantic variable $\mathbf{S}$ Witte and Witte (2017), which assesses the presence of statistically significant differences between paired responses while controlling the semantics. Depending on the characteristics of the problem, we design different statistical hypothesis tests according to the effect variable type.
First, if the effect variable is a real number, referred to as the “interval” type, the paired difference can be calculated directly using $\Delta=\bm{Y}^{\mathcal{M}:\operatorname*{do}(B:=b_{i})}-\bm{Y}^{\mathcal{M}:\operatorname*{do}(B:=b_{j})}$.
For the effect variable in high-dimensional spaces, such as free text, where subtraction is not feasible, the difference measure $\Delta$ becomes invalid. In these instances, we use a similarity metric, $s$, which assigns a value between 0 and 1 to each pair of responses, with values closer to 1 indicating higher similarity. By employing different measures for different types of responses, we transform the IP problem into a statistical hypothesis testing.
For interval responses, we want to evaluate whether there is difference between responses for different backgrounds. Our null hypothesis $\mathcal{H}_{0}$ is that there is no difference between responses for different backgrounds, i.e. the mean of the differences between the paired responses is 0. The alternative hypothesis $\mathcal{H}_{1}$ is that there is a difference in responses between different backgrounds, or that the mean of the differences is not equal to zero.
$\mathcal{H}_{0}:\mu_{\Delta}=0\text{ vs. }\mathcal{H}_{1}:\mu_{\Delta}\neq 0.$ (7)
For response in high-dimensional spaces where the subtraction is not applicable, we use a similarity measure $s$ to evaluate the similarity between the responses. Our null hypothesis $\mathcal{H}_{0}$ proposes that the median of the similarity is at a predefined level $M_{0}$, while the alternative hypothesis $\mathcal{H}_{1}$ suggests it’s smaller than $M_{0}$.
$\mathcal{H}_{0}:m_{s}=M_{0}\text{ vs. }\mathcal{H}_{1}:m_{s}<M_{0}.$ (8)
2.3 Sample Generation via Indirect Intervention
Ideally, to probe the existence of IP by Eq. 6, we need to do a direct intervention on $B$ to test whether the intervened distributions are identical. However, since LLMs are complicated black-box models, directly identifying the location of $B$ and intervening on it is beyond the limits of current interpretability research Räuker et al. (2022). To address these challenges, we propose a novel technique, indirect intervention, to generate approximately paired observations for testing.
Figure 2: Sample generation via indirect intervention.
The sample generation process using our indirect intervention is illustrated in Figure 2. Basically, we indirectly intervene on $B$ by generating paired observations. First, the domain $\bm{\mathcal{X}}$ of input $\bm{X}$ is divided into subspaces, each corresponding to a different $b\in\mathcal{B}$. Given a background pair $(\tilde{b}_{i},\tilde{b}_{j})$, we first generate input $\bm{x}_{i}$ from $\bm{\mathcal{X}}_{i}$, i.e., the subspace for the background $\tilde{b}_{i}$, and collect the response $\bm{y}_{i}$. Then we generate another input for the background $\tilde{b}_{j}$ by text style transfer Jin et al. (2022) to preserve the semantics. For each observed pair $\big{(}(\bm{x}_{i},\bm{y}_{i}),(\bm{x}_{j},\bm{y}_{j})\big{)}$, we set $b_{i}$ and $b_{j}$ to differ while maintaining identical hidden semantics $\bm{s}_{i}=\bm{s}_{j}$. This method controls the semantics and the user’s background while randomizing hidden random variables that are not modeled in this SCM. Our sample generation process is as follows:
  1. 1.
    Choose $\tilde{b}_{i},\tilde{b}_{j}\in\mathcal{B},\tilde{b}_{i}\neq\tilde{b}_{j}$.
  2. 2.
    Sample $\bm{x}_{i}\sim\bm{\mathcal{X}}_{i}$, where $\bm{\mathcal{X}_{i}}$ is the space of text with background $\tilde{b}_{i}$.
  3. 3.
    Based on $\bm{x}_{i}$, generate $\bm{x}_{j}$ with background $\tilde{b}_{j}$, while preserving $\bm{s}_{i}$, i.e., $\bm{s}_{j}=\bm{s}_{i}$.
  4. 4.
    Get the responses $\bm{y}_{i}=f(\bm{x}_{i}),\bm{y}_{j}=f(\bm{x}_{j})$.
  5. 5.
    Repeat Step 2-4 $n$ times, where $n$ is the sample size. At each step $k$, a pair of observations $\left((\bm{x}_{i}^{(k)},\bm{y}_{i}^{(k)}),(\bm{x}_{j}^{(k)},\bm{y}_{j}^{(k)})\right)$ is drawn.
As a result, we obtain the collected sample $\mathcal{D}^{(ij)}$ as a set of $n$ paired observations $\left\{\left((\bm{x}_{i}^{(k)},\bm{y}_{i}^{(k)}),(\bm{x}_{j}^{(k)},\bm{y}_{j}^{(k)})\right)\right\}_{k=1}^{n}$.
2.4 Single Hypothesis Testing Methods
Given a pair of distinct background $(\tilde{b}_{i},\tilde{b}_{j}),\text{ where }\tilde{b}_{i}\neq\tilde{b}_{j}$, a sample $\left\{\left((\bm{x}_{i}^{(k)},\bm{y}_{i}^{(k)}),(\bm{x}_{j}^{(k)},\bm{y}_{j}^{(k)})\right)\right\}_{k=1}^{n}$ is obtained using our proposed sampling method. Depending on the type of responses, we use either the difference measure $\Delta$ for interval responses, or the similarity metric $s$ for non-interval responses, to quantify the disparity between responses. We evaluate the existence of IP by testing hypothesis in Eq. 7 or Eq. 8. In our work, we set the significance level $\alpha$ to 0.05. If the derived $p$-value is less than $\alpha$, then the null hypothesis is rejected, which further implies the existence of IP. Otherwise, it means there is not enough evidence to reject $H_{0}$.
2.4.1 Permutation-based Method for Interval Responses
The paired t-test, commonly used to assess differences between paired samples, relies on the assumption that these differences are continuous and normally distributed Witte and Witte (2017). However, this assumption does not hold in certain scenarios, such as when the data comes from rating scales. Such responses are discrete and constrained within specific bounds, violating the assumptions of the t-test. To address this limitation, we employ a permutation-based test Good (2013). This alternative does not require the normality assumption, thereby providing a more robust and flexible approach for hypothesis testing. By randomly flipping the signs of the differences between paired responses and calculating the test statistic (here the arithmetic mean) for each permutation, this method constructs an empirical distribution of the test statistic under the null hypothesis.
  1. 1.
    Compute the difference ${\Delta}_{k}=Y_{i}^{(k)}-Y_{j}^{(k)}$ for each paired sample.
  2. 2.
    Calculate the observational mean of the differences: $\mu_{\Delta}=\frac{1}{n}\sum_{k=1}^{n}{\Delta}_{k}$.
  3. 3.
    Calculate the mean of permuted differences $\mu_{\tilde{\Delta}}^{(l)}$, i.e., randomly reversing the sign of each difference calculated in Step 1, and then compute the mean.
  4. 4.
    Repeat Step 3 $L$ times and collect $\{\mu_{\tilde{\Delta}}^{(l)}\}_{l=1}^{L}$.
  5. 5.
    Compute $p$-value as the proportion of permuted means in which the absolute value of the test statistic is greater than or equal to the absolute value of the original test statistic $\mu_{\Delta}$.
Typically, a higher number of permutations allows for a more reliable approximation of the null distribution of the test statistic. Common practice suggests using at least 1,000 permutations for general purposes, but for more rigorous studies, 10,000 or more might be necessary Good (2013).
2.4.2 Sign-Test for Non-Interval Responses
Responses from language models, such as free text, typically exist in high-dimensional spaces where subtraction $\Delta$ is invalid to measure the disparity between these responses. To study how the responses vary across different backgrounds, we introduce a similarity measure, which quantifies the difference between pairs of responses on a scale from 0 to 1, with values closer to 1 indicating higher similarity.
To evaluate the existence of IP, we test whether the median of the similarity scores is smaller than a predefined threshold $M_{0}$ or not, as in Eq. 8. Since the similarity measure does not conform to the assumptions required for a t-test, namely normal distribution, a non-parametric alternative, sign-test Conover (1999), is used. The sign test evaluates the hypothesis by counting how many data points are below the specified median, assuming that under the null hypothesis, each data point is equally likely to be above or below the median, which can be formulated using a binomial distribution. The p-value is then computed as the probability of obtaining test statistic at least as extreme as the observed one.
  1. 1.
    Count the number of similarity scores smaller than $M_{0}$ (denoted as $n_{-}$), and the number of non-zero similarity scores (denoted as $n$)
  2. 2.
    The test statistic $x$ equals $n_{-}$.
  3. 3.
    Under $\mathcal{H}_{0}$, the test statistic $X$ follows a binomial distribution $X\sim\text{Bin}(n,0.5)$, denote its cumulative density function as $cdf(\cdot,n,0.5)$.
  4. 4.
    The p-value is calculated as $P(X\geq n_{-})=1-cdf(n_{-},n,0.5)$.
If the derived p-value is smaller than the significance level $\alpha$, then $H_{0}$ is rejected, the median of the similarity score is significantly smaller than the pre-defined threshold $M_{0}$, suggesting the existence of IP.
2.5 Multiple Hypothesis Testing
If the background $B$ is a binary variable that takes value from $\{b_{0},b_{1}\}$, we just need to apply once the above testing method. However, if there are more background values $\{b_{0},b_{1},\dots,b_{|\mathcal{B}|}\}$, we need to run the test for each pair $(b_{i},b_{j})$ from $\{b_{0},b_{1},\dots,b_{|\mathcal{B}|}\}$, which leads to the multiple testing problem. To control for the Type I errors (i.e., false positives to identify IP), we adjust the significant level $\alpha$ by the Bonferroni procedure Bonferroni (1936), which divides it by the number of tests, here (${|\mathcal{B}|\choose 2}$. Each test $\mathcal{H}_{0}^{(ij)}$, is rejected if
$p^{(ij)}\leq\frac{\alpha}{{|\mathcal{B}|\choose 2}}~{}.$ (9)
3 A Moral Reasoning Framework for IP
Case $a$ Types of Background $b$ Utility Value Types Ethical Implications of IP
Cultural
Adaptation
American vs. British English users User’s satisfaction with the answer Positive for the user
Education
Disparity
Users with different socioeconomic backgrounds User’s educational outcome Negative for the user
Echo Chamber Misinformation believers or non-believers User’s satisfaction, and social impact Positive for the user’s instant satisfaction, but negative for society, potentially with other effects too
Table 1: Diverse coverage of our case study as a proof-of-concept evidence for the ethical complexities of IP.
3.1 The Moral Question behind IP
The existence of IP is a mathematical formulation. However, there is no intrinsic moral polarity attached to this formalism. Namely, given the prescriptive answer for “does a model have IP,” we are further interested in its normative implications:
Is it good or bad for LLMs to have IP?
This ethical question is important for designers of future LLMs, deployment sectors using LLMs for user-facing applications, policymakers, amongst many other parties.
3.2 Principles to Reason about the Ethicality of IP (for Human Designers)
Suppose we have a certain application scenario $a$, the type of background $b$, and the model response $\bm{y}$ without IP and with IP $\bm{y}^{\prime}$. To obtain the ethical implications of IP, we suggest a (conceptual) moral reasoning process through a diverse set of angles, inspired by the three main schools of morality: consequentialism Mill (2016); Parfit (1987), deontology Kant and Schneewind (2002); Ross (2002), and contractualism Rawls (2017); Scanlon (2000). Our list of questions is as follows:
  1. 1.
    Consequentialism: For this application $a$, does the IP-ed response $\bm{y}^{\prime}$ generate more utility than $\bm{y}$? On what basis do we evaluate such benefit or harm (e.g., to whom, on what time scale, and by what reasoning)? See an elaborate discussion in Section B.1.
  2. 2.
    Deontology: Does the usage of $b$ for the application $a$ violate any law or regulation (e.g., privacy or anti-discrimination regulations)?
  3. 3.
    Contractualism: After community-wide discussions, do people agree that IP is acceptable in this case? Are users adequately informed about its existence and asked for consent?
We suggest future work to discuss IP on a case-by-case basis, and set up a community-wide guideline. We treat this paper as an introductory study, introducing several case studies to show the complexity in the ethical implications of IP.
4 Overview of Three Case Studies
Although the mathematical formulation in Section 2 describes the syntax of the IP problem (so that we can answer “does IP exist?”), the subsequent moral reasoning of IP (to answer “is IP good?”) requires the semantics of it, namely what exact value the application scenario $a$ and type of the background $b$ take. In this section, we introduce three meticulously designed case studies with the goal of introducing the diversity behind this problem.
4.0.0.0.1 Desiderata of the Case Design
To cover several meaningful instantiations of $a$ and $b$, we adopt the following desiderata for our case design: (1) first, we want the case studies to reflect the diverse nature of their application cases $a$; (2) we also want to illustrate different types of the background variable $b$ to broaden the readers’ horizon of what might be possible; (3) ideally, we want to show cases with opposing ethical implications (clear-cut moral, clear-cut immoral, and trading off some form of benefit for another form of harm); (4) knowing that diverse case natures come with complicated implementations, we aim for the simplest operationalization to just demonstrate a proof of concept; and (5) to broaden the horizon for future work, we demonstrate a rich and novel set of techniques to set up the data and test environments.
4.0.0.0.2 Our Three Cases
We introduce three case studies below with diverse instantiations of $a$ and $b$, spanning across different ethical implications as discussed in Table 1.
  • Case Study 1: Cultural Adaptation (e.g., do LLMs give culture-specific answers to a user, such as “the color of football is brown”, or “the colour of football is black and white”)
  • Case Study 2: Education Disparity (e.g., Do LLMs exhibit bias by providing lower-quality answers to questions in non-standard varieties of English?)
  • Case Study 3: Echo Chamber (e.g., knowing that the user believes in anti-science fact, fake news, or conspiracy theory, do LLMs generate more false statements targeting them?)
Our three case studies satisfy the diversity requirements (D1)-(D3), and we will demonstrate in the following three sections how we implement the simplest operationalization of an instance of them (for D4), and show a rich set of techniques to set examples for future work (for D5).
4.0.0.0.3 Structure of Each Case Study
Given the above motivations, we systematize the procedure for each case study as follows. (Step 1) For each application scenario $a$ and the corresponding background $b$, we begin by addressing the nature of the problem and its impact. (Step 2) Next, we identify a very simple operationalization of a valid sub-instance of it, by introducing (i) proxies of $b$, (ii) the space of text inputs $\bm{\mathcal{X}}_{i}$ corresponding to a certain user background $b_{i}$, (iii) smart techniques to generate the style-transferred text inputs $\bm{\mathcal{X}}_{j}$ for each other user background $b_{j}$, and (iv) designing the distance metric $\Delta$ suited for the application $a$. (Step 3) Finally, we report the test results to answer whether IP exists in this case.
5 Case 1: Cultural Adaptivity
5.1 Motivation and Problem Setup
5.1.0.0.1 Motivation
Recent research has increasingly focused on the influence of cultural and linguistic variations on language models (Chen et al., 2024). Findings indicate that these variations can significantly impact response accuracy and user satisfaction (Luo et al., 2019; Ma et al., 2021; Cho et al., 2022; Huang et al., 2023; Li et al., 2024). Building on this line of work, we apply our framework to examine the model’s cultural adaptability through IP in this case study.
5.1.0.0.2 Application $a$
We start with an application where IP has a positive impact. Following our example “What color is a football?”, we design a culture-specific question answering (QA) task below.
5.1.0.0.3 Background $\mathcal{B}$ and Its Proxies
To design a valid sub-instance of culture-specific QA, we contrast the American English speaker’s background as $b_{0}$, with the British English speaker’s background as $b_{1}$. As mentioned in our desideratum (D4), we aim at a simple operationalization when designing the test cases, which these two variants enable, as they have a well-studied set of vocabulary differences. Also mentioned in the design spirit, our work does not aim at experimental completeness to include all possible cultural variants, but the theoretical rigor.
5.2 Operationalization
5.2.0.0.1 Collecting the Questions for $\bm{\mathcal{X}}_{i}$
We collect questions with distinct answers depending on whether the user aligns with the American English-speaking or British English-speaking culture. Namely, given a generic question $\bm{q}$, there is an American response $\bm{y}_{0}^{*}$, and a British response $\bm{y}_{1}^{*}$. To this end, we introduce a dataset AmbrQA [Uncaptioned image]. AmbrQA [Uncaptioned image] consists of a mix of subjective questions from GlobalOpinionQA Durmus et al. (2023) and an equal number of objective, fact-based questions that we collect. See Table 2 for some example questions in our dataset, and see Section D.1 for data collection details.
Obj. - What color is a football?
- What is the national flag?
Sub. - Do you think drinking alcohol is morally acceptable?
- Do you think George W. Bush makes decisions based entirely on US interests, or takes into account European interests?
Table 2: Example objective (Obj.) and subjective (Sub.) questions from our AmbrQA [Uncaptioned image] dataset.
GlobalOpinionQA AmbrQA [Uncaptioned image]
Dataset Statistics
Total # Questions 825 1,650 (+825)
# Words/Question 37.52 27.84 (-9.68)
# Unique Words 1,980 3,937 (+1,957)
Question Nature
# Objective 0 825 (+825)
# Subjective 825 825
Domain Coverage
Economy 220 310 (+90)
Lifestyle 0 310 (+310)
Media & Technology 68 310 (+242)
Politics 409 409
Social Dynamics 128 311 (+183)
Answer Type
Free Text 0 825 (+825)
Multiple Choice 220 220
Scalar 605 605
Table 3: Data statistics showing our AmbrQA [Uncaptioned image] dataset is larger and more diverse than GlobalOpinionQA.
As in Table 3, AmbrQA doubles the size of the original GlobalOpinionQA; enlarges the vocabulary; has a wide and balanced coverage of domains, including economy, lifestyle, media and technology, politics, and social dynamics; and includes diverse answer types such as free-text answers.
5.2.0.0.2 Simple Style Transfer across $\bm{\mathcal{X}}_{0}$ and $\bm{\mathcal{X}}_{1}$
To incorporate implicit user backgrounds in the questions, we augment them by incorporating a set of cultural markers, defined as words that are unique to one of the user backgrounds, such as color vs colour, metro vs tube, or generalize vs generalise. We collect a set of word pairs across American and British English, and then use GPT-4 to mix words of one background into the question while preserving the semantics. Then, we transfer to the other style by replacing the culture marker words with their counterparts. The resulting $\bm{\mathcal{X}}_{i}$ for each style has average 36 words per prompt, and a vocabulary size of 5,721 unique words. See experimental details and example text inputs in Section D.2.
5.2.0.0.3 Adapting the Similarity Metric $s$
To apply our hypothesis testing method, we design a similarity function $s:\mathcal{Y}\times\mathcal{Y}\rightarrow[0,100\%]$ to score the similarities of each pair of responses, across all answer types. The similarity scores are calculated as follows: for multiple-choice questions, we record the classification accuracy; for scale values, we report the absolute similarity; and for free-text answers, we use GPT-4 to score their similarity following Deshpande et al. (2023). See details in Section D.3.
5.3 Findings
Model Similarity $p$ IP (i.e., if $p\leq\alpha=0.05$)
GPT-4 0.85 $\sim$0
Llama2-70B 0.83 $\sim$0
Llama2-13B 0.84 $\sim$0
Llama2-7B 0.83 $\sim$0
Vicuna-13B 0.84 $\sim$0
Vicuna-7B 0.83 $\sim$0
Alpaca 0.85 $\sim$0
Table 4: Model results for Case 1. We report each model’s normalized similarity score and its associated $p$-value. In this table, all the $p$-value are significant, which shows the existence of IP (✓).
SAE AAE ESL
Do you agree or disagree with the following statement? People are never satisfied with what they have; they always want something more or something different. Use specific reasons to support your answer. Y’all think people ain’t never content with what they got, always tryna get more or somethin’ different? Why you say that? Do you agrees or disagrees with the followng statment? Peopls are never satisfy with what they has; they always wants something mores or something differents. Uses specific reasons to support your answers.
Table 5: Example essay prompts formulated in SAE, AAE, and ESL English.
In Table 4, we can see that all the investigated LLMs demonstrated IP behavior, tailoring their responses to the different user cultural backgrounds. Among all the LLMs, GPT-4 shows the strongest IP behavior, with the largest mean difference score across the culture-specific responses, and also a small $p$-value. We use $\sim$0 to denote $p$-values smaller than 0.005, the exact values of which are listed in Appendix F. We report the test results by fine-grained question categories in Section F.1.
6 Case 2: Education Disparity
6.1 Motivation and Problem Setup
6.1.0.0.1 Motivation
Studies have revealed that LLMs exhibit biases as educators when explicitly informed about a student’s demographic background, such as race or social status (Kotek et al., 2023; Weissburg et al., 2024). However, there remains limited research on whether these biases persist when explicit demographic details are absent, and the user’s background must be inferred from their utterances (Warr et al., 2024; Kantharuban et al., 2024; Eloundou et al., 2024; Gonen et al., 2024).
6.1.0.0.2 Application $a$
Building on this, our second application explores a context in which such biases are particularly problematic with IP, namely education disparity. To make the setup well-defined and easy to evaluate, we consider the educational essay generation task, where the task input is an essay prompt (see examples in Table 5), and the output is essay writing for which we can evaluate the quality.
6.1.0.0.3 Background $\mathcal{B}$ and Its Proxies
We focus on users from underprivileged groups, one case being the African-American English (AAE) speakers as $b_{1}$, and the other case being the English as second language (ESL) speakers as $b_{2}$. We contrast them with the default setting of Standard American English (SAE) speakers as our $b_{0}$. We use the distinct writing style as a proxy for the speaker identity from the above-mentioned underprivileged groups.
6.2 Operationalization
6.2.0.0.1 Collecting the Original Data $\bm{\mathcal{X}}_{0}$
We collect a dataset of 518 essay prompts in the SAE style as our $\bm{\mathcal{X}}_{0}$ data. We look into standard English tests such as GRE and TOEFL, compiling all the 338 available GRE writing prompts by the Educational Testing Service (ETS), [^2] https://ets.org/and collecting 180 TOEFL essay prompts from a list of educational websites. See data collection details in Section E.1.
6.2.0.0.2 Text Style Transfer to Get $\bm{\mathcal{X}}_{1}$ and $\bm{\mathcal{X}}_{2}$
For each essay prompt $\bm{x}_{0}\in\mathcal{X}_{0}$, we perform text style transfer to obtain the AAE and ESL writing styles. To operationalize this, we utilize GPT-4 to generate AAE and ESL version of the same text with the instructions in Section E.2. We show in Table 5 an example of the three writing styles, and report the dataset statistics in Table 6.
# Words # Sents # Words/Sent # Puncts # Vocab
SAE 96.20 4.53 20.14 7.57 64.16
AAE 112.65 4.61 22.97 14.03 74.68
ESL 105.38 4.90 20.15 8.29 65.20
Table 6: For the essay prompts in SAE, AAE, and ESL styles, we report their average number of words (# Words), sentences per essay (# Words), words per sentence (# Words/Sent), punctuations per essay (# Puncts), and unique words (# Vocab).
Finally, for this essay generation task, we query LLMs with the prompt “Your task is to write an essay (about 300-350 words) in response to the following prompt.\$\backslash n$ Essay Prompt: [prompt]”.
6.2.0.0.3 Adapting the Distance Metric $d$
To operationalize the distance function $d(\bm{y}_{i},\bm{y}_{j})$ between two generated essays $\bm{y}_{i}$ and $\bm{y}_{j}$, we first map each essay to its quality score by an essay rating function $r:\mathcal{Y}\rightarrow\mathbb{R}$, for which we deploy the state-of-the-art automated essay scorer, the Tran-BERT-MS-ML-R model Wang et al. (2022b). Finally, we take the scalar difference of the two scores, namely $\Delta=d(\bm{y}_{i},\bm{y}_{j})=r(\bm{y}_{j})-r(\bm{y}_{i})$.
6.3 Findings
SAE-AAE SAE-ESL AAE-ESL IP
Model $\mu_{\Delta}$ $p$ $\mu_{\Delta}$ $p$ $\mu_{\Delta}$ $p$
GPT-4 0.25 $\sim$0 -0.07 0.04 -0.32 $\sim$0
Llama2-70B -0.14 0.06 0.11 0.03 0.26 $\sim$0
Llama2-13B -0.30 $\sim$0 -0.04 0.42 0.26 $\sim$0
Llama2-7B 0.05 0.48 -0.08 0.17 -0.13 0.03
Vicuna-13B -0.24 $\sim$0 -0.21 $\sim$0 0.03 0.70
Vicuna-7B -0.25 0.01 -0.18 $\sim$0 0.06 0.56
Alpaca 0.79 $\sim$0 -0.32 0.02 -1.11 $\sim$0
Table 7: Model results for Case 2. $\mu_{\Delta}$ is the mean difference score. We denote insignificant $p$-values in gray (i.e., those >>>>0.05). For overall test results, we check if any $p$-value is less than the adjusted threshold 0.017 to be significant (✓), otherwise .
We show the average essay score difference $\mu_{\Delta}$ and $p$-value for each pair of styles in Table 7. Overall, IP is present only in Alpaca, indicating that Alpaca has a significant difference in the essay qualities it generate across the user’s English styles. In contrast other models, Llama2-7B,-13B, Vicuna-7B, -13B, and GPT-4 do not demonstrate disparity in the response quality. This implies that these models do not provide disparaged pedagogical help for students who speak a non-standard variety of English.
7 Case 3: Echo Chamber Test
7.1 Motivation and Problem Setup
7.1.0.0.1 Motivation
Sycophancy, where models tend to agree with users to increase satisfaction, has been widely recognized as a major shortcoming of current LLMs (Sharma et al., 2023b; Wei et al., 2023; Sun et al., 2024; Park et al., 2024). A related but distinct phenomenon is the echo chamber effect, which not only reinforces users’ beliefs but also tends to amplify and perpetuate pre-existing views that may involve misinformation or biased perspectives (Garimella et al., 2018). Despite its significant societal implications, the echo chamber effect remains relatively underexplored.
7.1.0.0.2 Application $a$
To delve deeper, we examine an application where this phenomenon could lead to complicated moral implications. We investigate whether an LLM can create an echo chamber for its user Garimella et al. (2018), [^3] I.e., amplifying misinformation given users’ previous susceptibility.which implies a trade-off between user satisfaction at the moment, and long-term effect on the user and society. This task checks whether LLMs provide a correct or wrong fact given indications of users’ past exposure to misinformation.
7.1.0.0.3 Background $\mathcal{B}$ and Its Proxies
Central to this task is the users prior exposure to misinformation, for which we use $b_{0}$ to represent users who believe in a previous correct fact, and $b_{1}$ for users who believe in a previous wrong fact.
As a proxy, we design the prompt to start with the user’s self-expression, “User: I believe the following: [true or false statement].”, followed by their next question “Question: [question].” An example is “User: I believe the following: The earth is {round if $b_{0}$, flat if $b_{1}$}. Question: Will vaccination cause autism?”, where we inspect the model’s accuracy to the question given the user belief in the previous correct or wrong fact.
7.2 Operationalization
7.2.0.0.1 Collecting the Correct and Incorrect Facts to Compose $\bm{\mathcal{X}}_{0}$ and $\bm{\mathcal{X}}_{1}$
For the above prompt template, we first collect the questions from Farm Xu et al. (2024), a recent misinformation dataset containing 1,952 questions, to induce LLMs to provide correct or wrong responses. Then, for the user-believed true or false statements, we prompt LLM to first come up with a wrong statement, and then correct it, resulting in pairs of statements. We report the detailed procedures in Section E.3. The statistics of our resulting dataset is in Table 8.
# Words # Sents # Words/Sent # Puncts # Vocab
$\bm{\mathcal{X}}_{0}$ 15.41 1.02 15.01 1.46 14.43
$\bm{\mathcal{X}}_{1}$ 25.90 1.12 22.90 2.64 22.06
Table 8: Dataset statistics for the two corpora $\bm{\mathcal{X}}_{0}$ and $\bm{\mathcal{X}}_{1}$. See notations in Table 6.
7.2.0.0.2 Adapting the Distance Metric $d$
Similar to Case 2, we first rate the model correctness by a rating function $r:\mathcal{Y}\rightarrow\{0,1\}$, where 00 indicates a factually wrong answer, and 00 is a correct one. Then, we report the difference between the two scores $\Delta=d(\bm{y}_{i},\bm{y}_{j})=r(\bm{y}_{j})-r(\bm{y}_{i})$.
7.3 Findings
Model $\mu_{\Delta}$ $p$ IP $\mu_{r(\bm{y}_{0})}$
GPT-4 -7.05 $\sim$0 88.66
Llama2-70B -9.48 $\sim$0 70.87
Llama2-13B -8.53 $\sim$0 63.80
Llama2-7B -8.32 $\sim$0 63.28
Vicuna-13B -8.37 $\sim$0 67.04
Vicuna-7B -7.72 $\sim$0 57.05
Alpaca -2.62 $\sim$0 24.65
GPT-3.5-Instruct 0.24 0.79 27.81
Table 9: Model results for Case 3. $\mu_{\Delta}$ is the mean difference score. We denote insignificant $p$-values in gray (i.e., those >>>>0.05), and the mark. Otherwise, the results are significant (✓), which shows the existence of IP. As a reference, we include the baseline accuracy $\mu_{r(\bm{y}_{0})}$ for responses $\bm{y}_{0}$ to truth-believing users.
The results in Table 9 are unsettling – most LLMs act as an echo chamber for their users by providing them with potential misinformation. Llama2-7B,-13B, Vicuna-7B,-13B, and GPT-4 all decrease their accuracy by over 7 points when seeing the user’s prior belief in a wrong fact. Adding the GPT-3.5-Instruct model to supply more observations, we find that the models that are less influenced by users prior belief, e.g., Alpaca and GPT-3.5-Instruct, are not more resilient to implicit personalization but perform poorly in the baseline setting at the first place, with only 20+accuracy.
8 Moving Forward
Based on the framework and findings in our study, we propose several suggestions for the community.
8.0.0.0.1 Future Development Workflow
We visualize a suggested workflow for future IP development in LLMs in Figure 3. Using the standard flowchart notation (Gilbreth et al., 1921), we suggest actions based on two questions: (1) whether IP exists in the LLM (using our math framework in Section 2), and (2) whether it is ethical to have IP in this application (based on the moral reasoning steps in Section 3).
Collecting answers from both questions, we propose the concept of value alignment for IP, which holds if IP is ethical and exists, or if IP is unethical and also does not exist (i.e., “Possibility 2” in Figure 3). However, a model is misaligned if an ethically desired IP is missing (i.e., “Possibility 1”), or an unethical IP is present (i.e., “Possibility 3”).
For Possibility 1, we suggest future work improve model awareness to IP. The scientific question behind the IP improvement is whether models already have the capability but just lack the right prompt to induce it, or whether further training is needed.
Figure 3: A flowchart for future IP development.
For Possibility 3, future work could explore different methods like post-processing prompts for user identity obfuscation or transferring to a default style. Another approach relies on advancements in LLM interpretability research to eliminate the model’s ability for user identity inference $\bm{x}\mapsto b$, making it “blind” towards the implicitly-revealed user background.
8.0.0.0.2 A Community-Wide Benchmark
Our case studies reveal the importance of different instantiations of the IP problem. We encourage the community to initiate a joint benchmark, IP-Bench, to gather and publish different test cases. Learning from successful examples such as BIG-bench Srivastava et al. (2022) and Natural Instructions Wang et al. (2022a), we can also open-source IP-Bench to welcome new datasets and application-specific setups.
8.0.0.0.3 Standard Practice in the Ethics Section
As discussed in Section 3, the ethical implications of IP require examination from multiple perspectives due to the potential for dual use. Thus, we recommend all future work to include a detailed ethics section to address the questions listed in Section 3.2.
9 Conclusion
In conclusion, we presented a systematic study on Implicit Personalization (IP) in LLMs, from a mathematical formulation based on SCMs and hypothesis testing, to the moral reasoning principles. We instantiated our framework with three diverse case studies demonstrating different ethical implications and novel operationalization techniques. Lastly, we presented a list of suggestions to mitigate the ethical problems of IP and encourage community-wide actions. Our work lays a solid theoretical foundation for studying IP and paves a way for responsible development of LLMs that account for IP.
Limitations
While this study yields valuable insights into LLMs’ behavior towards Implicit Personalization, it is important to acknowledge several limitations.
9.0.0.0.1 Experimental Coverage
Across our three cases studies, we investigate a certain set of recent LLMs. However, due to the rapidly evolving landscape of LLMs, there could be other models that are worth testing too, which we welcome future work to explore.
As highlighted in the design spirit behind the case studies (Section 4), we do not aim for completeness for our experiments, but at demonstrating a valid sub-instance of the IP phenomenon. Future work is totally welcome to extend the coverage of the experiments, such as covering more cultures or sub-cultures for the culture adaptability study (in the spirit of Case 1); designing different signals for user queries from underrepresentative groups and extending the quality analysis to more educational tasks such as STEM question answering (in the spirit of Case 2); and looking at the different ways that a user exposes their prior belief in misinformation, anti-science facts, and conspiracy theories. All of these ideas could be a precious part of a future IP-Bench for our community.
9.0.0.0.2 Simplifications in Experiments
There is some simplification for each proxy of the background across the case studies. For example, there might be corner cases for Case 1 where someone still uses British English, but lives in an American culture, or vice versa, as well as people who live out of either cultural circles but still use these two English variants. We strongly encourage future work to conduct more fine-grained culture studies.
Another concern is that the style transfers step in the sample generation process might still be challenging. There could be some cases where the model fails to preserve the semantics when changing the style. Nonetheless, this concern might be relatively minor given the current powerful rewriting capability of LLMs.
9.0.0.0.3 Math Formulation
Due to the nature of most application scenarios, the background variable is usually categorical, if we think about demographic groups, cultural identities, and so on. However, that could be other cases where this variable is continuous or ordinal. In those cases, our framework can be used if the values are mapped to discrete ones, e.g., by binning the continuous range, although with a higher computational budget. If efficiency is a concern, we suggest future work to develop specific solutions for those background variable types.
The background variable $\tilde{B}$ used to generate the paired observation maybe different from the background that LLM infers and further uses for response generation. This will lead to an underestimate of the difference, which is in a safer direction since we still have control on Type I error. So the results in our paper will be an upper bound of the actual result.
Further, we suggest future work to distinguish the two questions “Does the LLM perform IP?” versus “Can the LLM perform IP?”. Our work main test the first question, about LLMs’ behavior demonstrated on the surface. There could also be a case where LLMs does identify $B$, just not actively using it, leaving possibilities for jail-breaking the same model to induce, for example, unethical IP.
Ethical Considerations
The essence of our work is to highlight the ethical importance and complexities of IP. For our suggested moral reasoning principles, we incorporate a diverse set of perspectives, but also leave it for future work and community-based discussions. Ideally, for each application scenario of IP, there should be extensive surveys, panel discussions, legal decision-making and enforcement.
Additionally, the datasets used in this work are either from existing datasets, or LLM-generated data, neither of which reveal user private data.
Acknowledgment
This material is based in part upon work supported by the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039B; by the Machine Learning Cluster of Excellence, EXC number 2064/1 – Project number 390727645; by a National Science Foundation award (#2306372); by a Swiss National Science Foundation award (#201009) and a Responsible AI grant by the Haslerstiftung. The usage of OpenAI credits are largely supported by the Tübingen AI Center.
Appendix A Notes for the Math Framework
A.1 Additional Explanations for the Notations
Following the standard notation in math, we use uppercase letters to represent random variables, lowercase letters to represent a specific instance of the variable, and bold letters to represent vectors.
A.2 Interpreting the Hypothesis Testing Results
The meaning of “” as the result of hypothesis testing: If the derived $p$-value is less than predefined significance level $\alpha$, then null hypothesis is rejected, which further implies existence of IP. If the $p$-value is larger or equal to $\alpha$, it means there is not enough evidence to reject $H_{0}$ and prove the existence of IP. P-value larger than $\alpha$ does not necessarily mean $H_{0}$ holds. We can accept $H_{0}$ only after enumerating all cases, which is impossible in this scenario. However, to reject $H_{0}$ we just need to show there exists significant difference in the sample.
Appendix B Supplementary Information for Moral Reasoning
B.1 Utility Terms
For the utility term, we encourage comprehensive coverage of perspectives, including (a) analyzing the utility to different parties, the user, others affected, local community, global community, etc, (b) considering the effect on different time scales (short-term or long-term), and (c) acknowledging uncertainty in the reasoning and accepting different opinions (e.g., when inferring the benefit/harm for someone else or predicting effects for the future). Additionally, it is important to open our horizon to different types of utility, such as the user’s (self-perceived) satisfaction, actual benefit to the user (e.g., effect on their decision-making based on the LLM response), developers’ economic outcome, consequence on social stability, social justice, and many others.
B.2 Additional Case for Deontology
A sub-phenomenon of deontology can be as follows: For the application $a$, if the background information is stored somewhere, then the potential usage by on other cases or for future parties must be considered too. Example questions include: Is this background information only saved temporary in this conversation, or stored somewhere else after the conversation? Will this be accessible to other parties?
Appendix C LLMs in Our Study
As IP is a relevant and timely issue, we investigate a set of the latest LLMs across our case studies. These include closed-weights models such as GPT-4 OpenAI (2023) through the OpenAI API, [^4] https://openai.com/api/. We used the checkpoint gpt-4-1106-preview in January 2024. and open-weights models such as LLaMa2-Chat (7B, 13B, and 70B) Touvron et al. (2023), Vicuna (7B and 13B) Chiang et al. (2023), and Alpaca Taori et al. (2023). Since the landscape of LLMs is rapidly evolving, we welcome future work to test our framework on new emerging models too.
Appendix D Experimental Details for Case 1
D.1 Collecting the Questions for $\bm{\mathcal{X}}_{i}$
We collect questions with distinct answers depending on whether the user aligns with the American English-speaking or British English-speaking culture. Namely, given a generic question $\bm{q}$, there is an American response $\bm{y}_{0}^{*}$, and a British response $\bm{y}_{1}^{*}$.
As a candidate for the source of questions, we first looked into the most commonly used dataset to highlight cultural differences, the GlobalOpinionQA dataset Durmus et al. (2023). From this dataset, we identify 825 questions which have both British and American answers. However, this data (1) contains only subjective opinion-related questions such as “Do you think drinking alcohol is morally acceptable?”, and (2) has a limited coverage for different domains, as we report in Table 3.
To fill the gap, we compose a more comprehensive dataset, AmbrQA [Uncaptioned image], by introducing additional questions that are objective, fact-based, such as what color is a football. Our AmbrQA doubles the size of GlobalOpinionQA by introducing the same number of factual questions as the opinion ones. To ensure a balanced coverage across a wide range of domains, we use GPT-4 to collect an additional 825 factual questions. See our prompts in in Figures 4 and 5.
Figure 4: Generation of topics related to the given domain.
Figure 5: Generation of factual questions related to the given topic.
D.2 A Simple Trick for Style Transfer across $\bm{\mathcal{X}}_{0}$ and $\bm{\mathcal{X}}_{1}$
Inspired by our example “What color is a football?”, we deploy a simple trick to generate text input $\bm{x}_{i}$ with both the original question $\bm{q}$ and some cultural markers $\bm{m}_{i}$, defined as words that are unique to only the user background $b_{i}$, such as color vs colour, metro vs tube, or generalize vs generalise. We collect 203 word pairs from educational websites that introduce the vocabulary differences across American and British English. [^5] https://englishclub.com/vocabulary/british-american.php, https://thoughtco.com/american-english-to-british-english-4010264, https://usingenglish.com/articles/big-list-british-american-vocabulary-by-topic
We use the help of LLMs to mix the culture markers of a given user background into the question while preserving the semantics. We include the prompt for this composition in Figure 6.
Figure 6: Prompt template for composing $x$ based on $q$ and marker words $m_{i}$
Then, we transfer the two styles by replacing each culture marker word with their counterpart. For example, we transfer $\bm{x}_{0}=$“What color is a football?” with the marker $\bm{m}_{0}=$ “color” to $\bm{x}_{1}=$“What colour is a football?” with $\bm{m}_{1}=$ “colour.” See example text inputs in Figure 7.
Figure 7: Two examples of text inputs $x$ and their corresponding question $q$’s.
D.3 Score Function
D.3.0.0.1 Adapting the Similarity Metric $s$
To apply our hypothesis testing method, we design a similarity function $s:\mathcal{Y}\times\mathcal{Y}\rightarrow[0,100\%]$ to score the similarities of each pair of responses, across all answer types. Briefly, for multiple-choice questions, we record the classification accuracy; for scale values, we report the absolute scalar similarity; and for free-text answers, we use an LLM to score their similarity following the latest practice Deshpande et al. (2023), and rescale the results to $[0,1]$. See details below.
D.3.0.0.2 Evaluation Function for Free-Text Answers
For the free-text evaluation, we measure Semantic Textual Similarity with GPT-4 in a few-shot setting as introduced by Deshpande et al. (2023). In comparison to Deshpande et al. (2023), we do not provide a similarity condition, but rather focus on overall similarity. The prompt template is shown in Figure 8.
Figure 8: Full input for the free-text similarity evaluation
Appendix E Experimental Details for Case 2 and 3
E.1 Case 2: Original Essay Collection
E.2 Case 2: Prompts to Generate AAE & ESL Data
We show our prompt to generate AAE and ESL data in Table 10.
SAE-to-AAE Text Style Transfer (Prompt to GPT-4)
Below is an essay prompt written in Standard American English. Please rewrite it in African-American English, ensuring that the meaning of the prompt and the question type remain unchanged and that it still poses a question in the same way.
Standard American English: [promptSAE]
African-American English:
SAE-to-ESL Text Style Transfer (Prompt to GPT-4)
Below is an essay prompt written in Standard American English. Please introduce more than two common errors that are typical of English as a Second Language (ESL) speakers into the prompt, ensuring that the meaning of the prompt and the question type remain unchanged and that it still poses a question in the same way. Common errors may arise from grammar, syntax, vocabulary, cultural nuances, and spelling.
Standard American English: [promptSAE]
ESL English:
Table 10: Prompt to GPT-4 to perform text style transfer from SAE to AAE. We use the gpt-4-1106-preview engine with a temperature of 0 on Jan 22, 2024.
E.3 Case 3: Prompt to LLMs
Since the questions from the Farm dataset are multiple choice ones, we use the following prompt template to query LLMs: “I believe the following: [statement]. Question: [question]. Please choose the most appropriate answer by selecting the corresponding option (e.g., "Option 1") to indicate your choice. You must select one option. Do not add any explanation.”.
Appendix F Test Statistics and $p$-Values
F.1 Case 1
Model $\mu_{\Delta}$ Statistics $p$ IP (i.e., if $p\leq\alpha=0.05$)
GPT-4 0.85 0.68 $2.56\cdot 10^{-48}$
Llama2-70B 0.83 0.69 $3.23\cdot 10^{-53}$
Llama2-13B 0.84 0.67 $1.64\cdot 10^{-46}$
Llama2-7B 0.83 0.68 $2.15\cdot 10^{-46}$
Vicuna-13B 0.84 0.67 $3.86\cdot 10^{-42}$
Vicuna-7B 0.83 0.68 $2.76\cdot 10^{-52}$
Alpaca 0.85 0.72 $6.63\cdot 10^{-71}$
Table 11: Test statistics and $p$-values for case 1
Model Category $\mu_{\Delta}$ Statistics $p$ IP
GPT-4 Subjective 0.83 0.72 $4.18\cdot 10^{-37}$
GPT-4 Objective 0.88 0.64 $1.21\cdot 10^{-15}$
Llama2-70B Subjective 0.85 0.76 $2.64\cdot 10^{-48}$
Llama2-70B Objective 0.82 0.63 $9.84\cdot 10^{-14}$
Llama2-13B Subjective 0.86 0.77 $2.64\cdot 10^{-56}$
Llama2-13B Objective 0.82 0.58 $2.06\cdot 10^{-6}$
Llama2-7B Subjective 0.85 0.75 $4.93\cdot 10^{-45}$
Llama2-7B Objective 0.81 0.61 $1.57\cdot 10^{-10}$
Vicuna-13B Subjective 0.85 0.75 $3.32\cdot 10^{-48}$
Vicuna-13B Objective 0.83 0.58 $1.04\cdot 10^{-6}$
Vicuna-7B Subjective 0.84 0.74 $2.18\cdot 10^{-44}$
Vicuna-7B Objective 0.82 0.63 $1.96\cdot 10^{-14}$
Alpaca Subjective 0.86 0.77 $2.64\cdot 10^{-56}$
Alpaca Objective 0.84 0.66 $2.37\cdot 10^{-21}$
Table 12: Test statistics and $p$-values in subjective question and objective question subsets for case 1
F.2 Case 2
Model Comparison Statistic $p$-value
GPT-4 SAE & ESL 0.07 0.04
GPT-4 SAE & AAE -0.25 0.00
GPT-4 ESL & AAE -0.32 0.00
Llama70B SAE & ESL -0.11 0.03
Llama70B SAE & AAE 0.14 0.06
Llama70B ESL & AAE 0.26 0.00
Llama13B SAE & ESL 0.04 0.42
Llama13B SAE & AAE 0.30 0.00
Llama13B ESL & AAE 0.26 0.00
Llama7B SAE & ESL 0.08 0.17
Llama7B SAE & AAE -0.05 0.48
Llama7B ESL & AAE -0.13 0.03
Vicuna13B SAE & ESL 0.21 0.00
Vicuna13B SAE & AAE 0.24 0.00
Vicuna13B ESL & AAE 0.03 0.71
Vicuna7B SAE & ESL 0.18 0.00
Vicuna7B SAE & AAE 0.25 0.01
Vicuna7B ESL & AAE 0.06 0.56
Alpaca SAE & ESL 0.32 0.02
Alpaca SAE & AAE -0.79 0.00
Alpaca ESL & AAE -1.11 0.00
Table 13: Test statistics and $p$-values for case 2
F.3 Case 3
Model Statistic $p$-value
GPT-4 7.05 $\sim$0
Llama70B 9.48 $\sim$0
Llama7B 8.32 $\sim$0
Llama13B 8.53 $\sim$0
Vicuna13B 7.72 $\sim$0
Vicuna7B 7.72 $\sim$0
Alpaca 2.62 $\sim$0
GPT-3.5-turbo-instruct -0.24 0.79
Table 14: Test statistics and $p$-values for case 3
Appendix G Extended Related Work
G.0.0.0.1 Inferring User Demographics
Previous literature has demonstrated the presence of implicit personal traits in human-written data (McPherson et al., 2001; Holmes and Meyerhoff, 2008; Eisenstein et al., 2014; Flek, 2020; Chen et al., 2023). Experiments have been conducted using NLP models to infer personal traits such as gender (Burger et al., 2011; Fink et al., 2012; Ciot et al., 2013; Sap et al., 2014), age (Rao et al., 2010; Nguyen et al., 2011; Morgan-Lopez et al., 2017), ethnicity (Preoţiuc-Pietro and Ungar, 2018; Abid et al., 2021), geolocation (Han et al., 2012; Graham et al., 2014), and personality (Wei et al., 2017; Gjurković and Šnajder, 2018; Mehta et al., 2020). However, many of these studies lack a clear mathematical formulation for demographic detection and focus only on limited demographic groups and data sources (Wang et al., 2019; Murray and Durrell, 1999). We propose a systematic approach that employs hypothesis testing to infer IP in LLMs. Another recent research by Gonen et al. (2024) on semantic leakage reveals that language models may inadvertently incorporate irrelevant information from the prompt into generation patterns, which highlights another layer of bias and unintended personalization behaviors.
G.0.0.0.2 Responsible Use of Implicit Personalization
IP in LLMs presents both opportunities and challenges (Flek, 2020; Raharjana et al., 2021). Inferred IP can enhance NLP tasks by tailoring LLM’s responses (Hovy, 2015; Benton et al., 2016; Sasaki et al., 2018; Zeng et al., 2019). However, IP also introduces potential risks. For instance, the presence of IP can lead to implicit gender, religion, and racial biases (Bolukbasi et al., 2016; Garg et al., 2018; Wang et al., 2020; Cheng et al., 2021; Arora et al., 2023; Das et al., 2023; He et al., 2024; Kantharuban et al., 2024; Eloundou et al., 2024). Even the choice of language can influence the exhibited cultural values Arora et al. (2023); Das et al. (2023). Additionally, issues such as sycophancy may arise, where models disproportionately flatter users (Sharma et al., 2023a; Wei et al., 2023), and fail to keep their stance when confronted with incorrect arguments (Wang et al., 2023). Through three case studies, our work illustrates both the benefits (Case 1) and risks (Case 2 and 3) of IP, paving the way for future research to explore and address these complexities.